# HUME: Measuring the Human-Model Performance Gap in Text Embedding Tasks

### v2

@@ Abstract
Comparing human and model performance offers a valuable perspective for understanding the strengths and limitations of embedding models, highlighting where they succeed and where they fail to capture meaning and nuance. However, such comparisons are rarely made, as human performance on embedding tasks is difficult to measure. To fill this gap, we introduce HUME: Human Evaluation Framework for Text Embeddings. While frameworks like MTEB provide broad model evaluation, they lack reliable estimates of human performance, limiting the interpretability of model scores. We measure human performance across 16 MTEB datasets spanning reranking, classification, clustering, and semantic textual similarity across linguistically diverse high- and low-resource languages. Humans achieve an average performance of 77.6% compared to 80.1% for the best embedding model, although variation is substantial: models reach near-ceiling performance on some datasets while struggling on others, suggesting dataset issues and revealing shortcomings in low-resource languages. We provide human performance baselines, insight into task difficulty patterns, and an extensible evaluation framework that enables a more meaningful interpretation of the model and informs the development of both models and benchmarks. Our code, dataset, and leaderboard are publicly available at https://github.com/embeddings-benchmark/mteb.
Figure 1: Overall ranking of human performance versus 13 embedding models across 16 tasks. Human annotators achieve 4th place with a score of 77.6, demonstrating competitive but not dominant performance. The ranking reveals significant variation in model performance across different parameter scales and architectures. Darker shades of blue means a larger model.
@@ 1 Introduction
Embedding models are central to modern NLP systems, powering applications such as search, recommendation, semantic analysis, and information retrieval. Many benchmarks test the performance of embedding models, with the most comprehensive offering a diverse suite of tasks that test their generality and robustness (Muennighoff et al., 2022; Xiao et al., 2025; Enevoldsen et al., 2025). Despite these advances, the interpretation and quality of these scores are often unclear as it is absent of human performance references. Current metrics define performance in terms of theoretical maxima (e.g., MAP = 1.0) that assume perfect consensus on task outcomes. However, many NLP tasks inherently involve ambiguity and disagreement (Plank, 2022), making a model’s score difficult to meaningfully interpret without reasonable references. For instance, if a model achieves 0.85 MAP in reranking, it is unclear whether this should be considered strong, mediocre, or beyond what annotators typically achieve. This disconnect highlights the need for human-centered evaluation that contextualizes benchmark results.
To address this, we introduce HUME: a Human Evaluation framework for text embedding tasks. HUME evaluates annotator performance across four task categories: reranking, classification, clustering, and semantic textual similarity (STS), using 16 diverse datasets from the Massive Text Embedding Benchmark (MTEB) (Muennighoff et al., 2022), adapted for human annotation feasibility. Through multi-annotator experiments, we analyze task difficulty, quantify variation across humans, and compare results directly against state-of-the-art embedding models.
@@ 2.1 Text Embedding Models and MTEB
Text embedding models map natural language into dense vectors that capture semantic information. They have progressed from static embeddings (Mikolov et al., 2013; Pennington et al., 2014) to contextual encoders (Devlin et al., 2019; Liu et al., 2019) and more recently to models optimized for embeddings like Sentence-BERT (Reimers & Gurevych, 2019), E5 (Wang et al., 2022), and GTE (Li et al., 2023), powering applications such as search, classification, and clustering.
To evaluate these models across its diverse use-cases, MTEB (Muennighoff et al., 2022; Enevoldsen et al., 2025) has risen as the de facto benchmark framework for embeddings and consolidates evaluation across diverse tasks and datasets. Despite its breadth and community-driven extensions – spanning multilingual, multimodal, and domain-specific variants (Xiao et al., 2025; Kasmaee et al., 2025; Tang & Yang, 2025; Xiao et al., 2024; Ciancone et al., 2024; Enevoldsen et al., 2024; Zinvandi et al., 2025; Wehrli et al., 2024; Poświata et al., 2024; Snegirev et al., 2024), MTEB lacks human performance baselines, making it difficult to contextualize model achievements.
@@ 2.2 Human Evaluation in NLP
Human evaluation is well established in NLP, especially for generative tasks like machine translation (Graham et al., 2013), summarization (Fabbri et al., 2021), and dialogue (Gupta et al., 2019). In contrast, embedding-based tasks have relied almost exclusively on automated metrics, with little attention to human baselines.
In information retrieval, initiatives such as TREC (Voorhees & Tice, 2000) collect human relevance judgments, but these serve as gold standards rather than benchmarks of human performance under model metrics (e.g., nDCG, MRR). Similarly, GLUE, SuperGLUE (Wang et al., 2018; 2019), and MERA (Fenogenova et al., 2024) – a Russian GLUE-like benchmark – report human baselines, but mainly for classification and reasoning tasks. For embeddings, works like STS (Cer et al., 2017) report inter-annotator agreement, yet these are not converted into model-comparable performance scores. This leaves a gap: human performance on embedding benchmarks such as MTEB remains largely unquantified.
@@ 3.2 Tasks, Datasets, and Metrics
Our selection criteria ensures comprehensive coverage across multiple dimensions: (1) linguistic diversity: including both high-resource languages (English, Arabic, Russian) and lower-resource languages (Norwegian Bokmål, Danish)  [^1] With 0 being ”The Left-Behinds” and 5 being ”The Winners”, we cover eng: 5, ara: 5, rus: 4, dan: 3, nob: 1 according to the 0-5 scale by (Joshi et al., 2021). to test cross-lingual generalization, (2) domain variety: spanning news, social media, encyclopedic content, scientific literature, and forum discussions to capture real-world application diversity, (3) construction methods: including both curated human annotations and synthetic dataset creation to understand how dataset origin affects human-model alignment, (4) task relevance: using tasks from established benchmarks widely adopted in the embedding evaluation community, and (5) task complexity variation: ranging from binary classification to fine-grained similarity judgments. This systematic selection ensures our findings generalize across the diverse landscape of embedding applications while maintaining direct relevance to existing evaluation frameworks.
Each task category uses a primary evaluation metric to enable consistent human–model comparisons. Table 1 summarizes the datasets, their domains, and the primary metrics applied. Detailed task examples are provided in Appendix B.
Datasets Description Metrics
Reranking Core17, News21, Robust04 (Weller et al., 2024) Information retrieval benchmarks (news, documents) MAP, MRR@10, nDCG@10
WikiMulti (Enevoldsen et al., 2025) Wikipedia article reranking (eng, dan, nob)
Classification Emotion (Saravia et al., 2018) Emotion classification from social media text Accuracy, F1, Weighted F1
Tweet Senti (Barbieri et al., 2022) Sentiment analysis of tweets
Toxicity (cjadams et al., 2019) Toxic content detection
Multilingual Sentiment (Mollanorozy et al., 2023) Sentiment classification (ara, eng, nob, rus)
Clustering WikiCities (Foundation, ) Entity clustering from Wikipedia V-Measure, ARI, AMI
ArXiv (arXiv.org submitters, 2024) Academic paper topic clustering (derived labels)
Reddit (Geigle et al., 2021) Forum discussion topic clustering
SIB200 (Adelani et al., 2023) Multilingual sentence clustering (ara, dan, eng, rus)
STS STSBenchmark (May, 2021) General semantic similarity benchmark Spearman, Pearson
SICK-R (Marelli et al., 2014) Semantic relatedness and entailment
STS12 (Agirre et al., 2012b) Shared task semantic similarity
STS22 (Chen et al., 2022) Multilingual semantic similarity (ara, eng, rus)
Table 1: Complete list of 16 datasets and evaluation metrics used for human annotation. For the metrics we use MAP (Manning et al., 2008), MRR (Manning et al., 2008), nDCG (Järvelin & Kekäläinen, 2002), Accuracy/F1 (Sokolova & Lapalme, 2009), V-Measure (Rosenberg & Hirschberg, 2007), ARI (Hubert & Arabie, 1985), AMI (Vinh et al., 2010), Spearman/Pearson (Spearman, 2010) (Pearson, 1895), following the MTEB implementations.
@@ 3.3 Annotation Procedure
All annotations are conducted in Argilla (Argilla Project Contributors, 2025) using task-specific interfaces: binary relevance judgments for reranking, categorical labels for classification, free assignment of cluster IDs for clustering, and 0–5 similarity scores for STS. Sample sizes are balanced by task complexity: reranking uses 20–49 queries per dataset, classification 40–48 examples, clustering 30 items across clusters, and STS 30–50 sentence pairs.
All annotators were male, aged 20–35, from culturally diverse backgrounds, and experienced NLP practitioners with native or near-native proficiency in the evaluated languages. They followed structured guidelines and completed all annotations independently, without access to ground truth or model predictions. The downsampled task subsets used for comparisons are included in the MTEB package, with detailed task examples provided in Appendix B.
English tasks were annotated by two annotators to enable agreement analysis. Multilingual tasks were annotated by a single annotator with corresponding language expertise. Inter-annotator agreement was assessed with task-appropriate metrics: Fleiss’ kappa (Fleiss, 1971) for classification, pairwise ARI (Strehl & Ghosh, 2003) for clustering, pairwise Spearman correlation (Agirre et al., 2012a) for STS, and mean Spearman/Kendall’s tau (Manning et al., 2008) for reranking. A detailed agreement analysis is provided in Appendix D. This controlled evaluation setup minimizes potential confounds from dataset variation and enables direct performance comparisons on identical evaluation instances.
@@ 3.4 Model Selection and Evaluation
We evaluate 13 embedding models chosen to cover multiple dimensions: (1) parameter scale (22M–7B), (2) architecture (encoder- and decoder-based), (3) instruction tuning (instruction-tuned and standard), and (4) multilingual capability (English and multilingual). This selection spans diverse computational budgets and training paradigms, capturing the current embedding landscape. All evaluated models are provided in Appendix F.
All models are evaluated on the downsampled instances annotated by humans, using identical metrics, protocols, and computational settings. Human performance is computed using the metrics in Table 1, mirroring MTEB protocols. For primary analyses, we report MAP for reranking, Accuracy for classification, V-Measure for clustering, and Spearman correlation for STS.
@@ 4 Results and Analysis
Figure 1 provides an overview of human performance relative to 13 state-of-the-art embedding models across 16 tasks. Human annotators rank 4th overall with a score of 0.776, trailing 3 large models but outperforming 10 others. This positioning reveals that humans neither represent a performance ceiling nor a lower bound, but rather occupy a middle ground that varies significantly by task category and language.
Table 2 presents the aggregated human results compared to selected models in various sizes. Below we analyze each category and language group in the following sections and refer to Appendix A for the full results on specific task categories.
@@ 4.1 Performance Patterns by Task Category
Figure 2 shows human performance relative to model performance ranges across all 26 task-language pairs. Each task shows human performance (point) positioned within the full spectrum from worst to best model performance (range bars). Humans consistently perform in the upper portion of model ranges, typically exceeding median model performance (61.5% of tasks) while rarely matching the best models (15.4% of tasks). Classification tasks show the strongest human performance, with humans outperforming all models in 3 of 7 tasks, while clustering and reranking reveal consistent gaps where humans fall short of top-performing models. Detailed gap analysis can be found in Appendix E.
Classification: Human performance averages 70.3 across classification tasks, with substantial variation reflecting task-specific challenges. Performance ranges from 45.8 on emotion classification ($\kappa=0.39$, fair agreement) to 95.0 on Arabic sentiment analysis. Models generally exceed human performance, with the best achieving 87.1 on English tasks. However, humans outperform models on non-English sentiment analysis, particularly in Arabic (95.0 vs. 77.5) and Russian (92.5 vs. 81.2), likely benefiting from native cultural and linguistic understanding that current models fail to capture.
Clustering
Clustering presents the greatest challenge for humans, averaging 67.4 V-measure with extreme variation. Near-perfect performance on WikiCities (97.6, $\text{ARI}=0.91$) contrasts sharply with poor performance on ArXiv papers (49.2, $\text{ARI}=-0.001$), which uses derived labels from ArXiv categories. Models consistently outperform humans, with the best reaching 85.1%. The poor inter-annotator agreement on academic clustering indicates fundamental task ambiguity rather than human limitation.
Reranking: Humans achieve strong performance in reranking (87.2 average MAP), demonstrating intuitive document relevance understanding. Models exceed human performance (best: 96.4), but humans maintain competitive scores with high inter-annotator agreement ($\rho=0.64$-$0.85$), suggesting these tasks align well with human judgment.
STS: Human performance averages 83.2 Spearman correlation, with notable dataset-specific variation. STS12 achieves 91.2 while STS22-Russian drops to 58.7, likely reflecting dataset quality issues discussed in § 4.3. Models achieve comparable performance (best: 88.5), with moderate inter-annotator agreement ($\rho=0.58$-$0.77$).

### v3

@@ Abstract
Comparing human and model performance offers a valuable perspective for understanding the strengths and limitations of embedding models, highlighting where they succeed and where they fail to capture meaning and nuance. However, such comparisons are rarely made, as human performance on embedding tasks is difficult to measure. To fill this gap, we introduce HUME: Human Evaluation Framework for Text Embeddings. While frameworks like MTEB provide broad model evaluation, they lack reliable estimates of human performance, limiting the interpretability of model scores. We measure human performance across 16 MTEB datasets spanning reranking, classification, clustering, and semantic textual similarity across linguistically diverse high- and low-resource languages. Humans achieve an average performance of 77.6% compared to 80.1% for the best embedding model, though with substantial variation: models reach high performance on some datasets while struggling on notably low-resource languages. Our human annotations also reveal multiple dataset issues. We additionally benchmark nine LLMs as annotators on reranking, classification, and STS tasks, finding that they fall short of human performance (76.1% vs. 81.2%) despite offering scalability advantages. We provide human performance baselines, insights into task difficulty patterns, and an extensible evaluation framework that enables a more meaningful interpretation of results and informs the development of both models and benchmarks. Our code, dataset, and leaderboard are publicly available at https://github.com/embeddings-benchmark/mteb.
Figure 1: Human performance versus 13 embedding models across 16 tasks. Humans rank 4th (77.6), showing competitive but not dominant performance. Darker shades indicate larger models.
@@ 1 Introduction
Embedding models are central to modern NLP systems, powering applications such as search, recommendation, semantic analysis, and information retrieval. Many benchmarks test the performance of embedding models, with the most comprehensive offering a diverse suite of tasks that test their generality and robustness (muennighoff2022mteb; xiao2025mieb; enevoldsen2025mmteb). Despite these advances, the interpretation and quality of these scores are often unclear as it is absent of human performance references. Current metrics define performance in terms of theoretical maxima (e.g., MAP = 1.0) that assume perfect consensus on task outcomes. However, many NLP tasks inherently involve ambiguity and disagreement (plank-2022-problem), making a model’s score difficult to meaningfully interpret without reasonable references. This interpretability gap has serious consequences. When benchmarks reward models for fitting noisy labels, labels where even human annotators disagree, the field risks "blind optimization": expending R&D resources to replicate annotation artifacts rather than achieve semantic progress. For instance, if a model achieves 0.85 MAP in reranking, it is unclear whether this should be considered strong, mediocre, or beyond what annotators typically achieve. This disconnect highlights the need for human-centered evaluation that contextualizes benchmark results. Importantly, human performance should not be treated as an upper bound but as a diagnostic signal: a way to understand where tasks are inherently noisy, where models may surpass typical annotator agreement, and where model behaviour diverges from human judgment.
To address this, we introduce HUME: a Human Evaluation framework for text embedding tasks. HUME evaluates annotator performance across four task categories: reranking, classification, clustering, and semantic textual similarity (STS), using 16 diverse datasets from the Massive Text Embedding Benchmark (MTEB) (muennighoff2022mteb), adapted for human annotation feasibility. Through multi-annotator experiments, we analyze task difficulty, quantify variation across humans, and compare results directly against state-of-the-art embedding models.
@@
Beyond human evaluation, we also investigate whether Large Language Models (LLMs) can serve as scalable proxies for human judgment. The promise of LLM-as-annotator approaches is compelling: if LLMs can reliably replicate human judgments, they could enable large-scale benchmark development at lower cost. However, this assumes LLMs capture the same semantic distinctions humans make, rather than exhibiting systematic biases. We evaluate nine state-of-the-art LLMs (GPT-5, GPT-4.1, Gemini, Mistral, and Qwen3 variants) on identical annotation tasks to assess their viability as human proxies and identify task-specific limitations.
@@ 2.1 Text Embedding Models and MTEB
Text embedding models map natural language into dense vectors that capture semantic information. They have progressed from static embeddings (mikolov2013distributed; pennington2014glove) to contextual encoders (devlin2019bert; liu2019roberta) and more recently to models optimized for embeddings like Sentence-BERT (reimers2019sentence), E5 (wang2022e5), and GTE (li2023gte), powering applications such as search, classification, and clustering.
To evaluate these models across its diverse use-cases, MTEB (muennighoff2022mteb; enevoldsen2025mmteb) has risen as the de facto benchmark framework for embeddings and consolidates evaluation across diverse tasks and datasets. Despite its breadth and community-driven extensions – spanning multilingual, multimodal, and domain-specific variants (xiao2025mieb; kasmaee2025chemteb; tang2025finmteb; xiao2024cpack; ciancone2024mtebfrench; enevoldsen2024scandinavian; zinvandi2025famteb; wehrli2024germantextembeddingclustering; poswiata2024plmteb; snegirev2024russianfocusedembeddersexplorationrumteb), MTEB lacks human performance baselines, making it difficult to contextualize model achievements.
@@ 2.2 Human Evaluation in NLP
Human evaluation is well established in NLP, especially for generative tasks like machine translation (graham2013continuous), summarization (fabbri2021summeval), and dialogue (gupta2019investigatingevaluationopendomaindialogue). In contrast, embedding-based tasks have relied almost exclusively on automated metrics, with little attention to human baselines.
In information retrieval, initiatives such as TREC (voorhees2000building) collect human relevance judgments, but these serve as gold standards rather than benchmarks of human performance under model metrics (e.g., nDCG, MRR). Similarly, GLUE, SuperGLUE (wang2018glue; wang2019superglue), and MERA (fenogenova2024meracomprehensivellmevaluation) – a Russian GLUE-like benchmark – report human baselines, but mainly for classification and reasoning tasks. For embeddings, works like STS (cer-etal-2017-semeval) report inter-annotator agreement, yet these are not converted into model-comparable performance scores. This leaves a gap: human performance on embedding benchmarks such as MTEB remains largely unquantified.
@@ 3.2 Tasks, Datasets, and Metrics
Our selection criteria ensures comprehensive coverage across multiple dimensions: (1) linguistic diversity: including both high-resource languages (English, Arabic, Russian) and lower-resource languages (Norwegian Bokmål, Danish)  [^1] With 0 being ”The Left-Behinds” and 5 being ”The Winners”, we cover eng: 5, ara: 5, rus: 4, dan: 3, nob: 1 according to the 0-5 scale by (joshi2021statefatelinguisticdiversity). to test cross-lingual generalization, (2) domain variety: spanning news, social media, encyclopedic content, scientific literature, and forum discussions to capture real-world application diversity, (3) construction methods: including both curated human annotations and synthetic dataset creation to understand how dataset origin affects human-model alignment, (4) task relevance: using tasks from established benchmarks widely adopted in the embedding evaluation community, and (5) task complexity variation: ranging from binary classification to fine-grained similarity judgments. This systematic selection ensures our findings generalize across the diverse landscape of embedding applications while maintaining direct relevance to existing evaluation frameworks.
Each task category uses a primary evaluation metric to enable consistent human–model comparisons. We summarize the datasets, their domains, and the primary metrics applied in Appendix A. Detailed task examples are provided in Appendix C.
Retrieval Proxy via Reranking
We use Reranking as a human-evaluable proxy for Information Retrieval. Direct human evaluation of large-scale retrieval is methodologically infeasible—requiring annotators to evaluate thousands of candidate documents per query. Reranking preserves the core semantic challenge of discriminating query-candidate relevance while remaining tractable: humans evaluate only the top-k candidates, establishing a baseline conceptually equivalent to embedding-space behavior and ensuring human-model comparability.
3.3 Instructions
Human instructions are designed to match the task definitions exactly (e.g., identical label sets for classification, same 1-5 scale for STS) to ensure valid comparisons. However, formal, detailed annotation protocols are not publicly available for many MTEB datasets, which limits our ability to verify alignment. To mitigate this, we designed instructions based on the original dataset papers’ task descriptions, ensuring annotators understood the semantic distinctions required for each task. Instructions were piloted with a small subset before full annotation to identify and resolve ambiguities.
@@ 3.4 Annotation Procedure
Our annotations process follows a trend similar to recent embedding benchmarks enevoldsen2025mmteb; xiao2025mieb focusing on a diverse set of tasks with fewer samples rather than large singular tasks. We choose this approach as it allow us to better cover the broad scopes of current benchmark. Annotators were recruited with a focus on cultural and language diversity.
All annotations are conducted in Argilla (argilla2025tool) using task-specific interfaces: binary relevance for reranking, categorical labels for classification, free cluster ID assignment for clustering, and 0–5 similarity scores for STS. Sample sizes balance task complexity: reranking (20–49 queries), classification (40–48 examples), clustering (30 items), and STS (30–50 pairs).
All annotators were male, aged 20–35, from culturally diverse backgrounds, and experienced NLP practitioners with native or near-native proficiency in the evaluated languages. They followed structured guidelines and completed all annotations independently, without access to ground truth or model predictions. The downsampled task subsets used for comparisons are included in the MTEB package, with detailed task examples provided in Appendix C.
English tasks were annotated by two annotators to enable agreement analysis. Multilingual tasks were annotated by a single annotator with corresponding language expertise. Inter-annotator agreement was assessed with task-appropriate metrics: Fleiss’ kappa (Fleiss1971) for classification, pairwise ARI (10.1162/153244303321897735) for clustering, pairwise Spearman correlation (agirre-etal-2012-semeval) for STS, and mean Spearman/Kendall’s tau (Manning_Raghavan_Schütze_2008) for reranking. A detailed agreement analysis is provided in Appendix F. This controlled evaluation setup minimizes potential confounds from dataset variation and enables direct performance comparisons on identical evaluation instances.
@@ 3.5 Model Selection and Evaluation
We evaluate 13 embedding models chosen to cover multiple dimensions: (1) parameter scale (22M–7B), (2) architecture (encoder- and decoder-based), (3) instruction tuning (instruction-tuned and standard), and (4) multilingual capability (English and multilingual). This selection spans diverse computational budgets and training paradigms, capturing the current embedding landscape. All evaluated models are provided in Appendix H.
All models are evaluated on the downsampled instances annotated by humans, using identical metrics, protocols, and computational settings. Human performance is computed using the metrics in Appendix A, mirroring MTEB protocols. For primary analyses, we report MAP for reranking, Accuracy for classification, V-Measure for clustering, and Spearman correlation for STS.
To account for sample size constraints, we determine statistical significance using 95% confidence intervals computed via Wilson Score Intervals (accuracy) and Fisher $z$-transformations (correlation), as detailed in Appendix E.
3.6 LLM-as-Annotator Evaluation
To assess whether automated evaluation can proxy human judgment, we evaluate nine state-of-the-art Large Language Models (LLMs) as annotators on the exact same tasks. We employ GPT-5 (full and mini), GPT-4.1 (full and mini), Gemini 2.5 Flash, Mistral Small-24B-Instruct, and three Qwen3 variants (30B, 32B, Coder-30B), prompting them with identical instructions provided to human annotators (see Appendix C).
LLMs receive the same task instances, evaluation metrics, and scoring protocols as human annotators, enabling direct performance comparisons. For classification and STS tasks, LLMs provide categorical labels or numerical similarity scores. For reranking, LLMs rank candidate documents by relevance to the query. Clustering tasks were excluded from LLM evaluation due to fundamental difficulties in eliciting consistent, structured cluster assignments from generative models.
This controlled setup determines whether LLMs can serve as scalable, low-cost proxies for human evaluation or whether they exhibit systematic biases that limit their utility for benchmark development. By evaluating LLMs on the same instances as humans and embedding models, we can directly compare their annotation quality and identify task-specific strengths and limitations.
@@ 4 Results and Analysis
Figure 1 provides an overview of human performance relative to 13 state-of-the-art embedding models across 16 tasks. Human annotators rank 4th overall with a score of 0.776, trailing 3 large models but outperforming 10 others. However, a raw ranking obscures the nuance of task difficulty and data reliability. As shown in Table 1, humans neither represent a uniform performance ceiling nor a lower bound, but rather occupy a middle ground that varies significantly by task category, language, and dataset quality.
We computed 95% confidence intervals for human performance using metric-appropriate methods (Wilson Score Intervals for classification accuracy, Fisher $z$-transformation for correlation-based metrics, and empirical annotator ranges for clustering and reranking). Models perform outside human CIs in 14 of 26 tasks ($p<0.05$), often on datasets with low inter-annotator agreement where “superhuman” performance may reflect artifact fitting rather than genuine capability (see Appendix E for methodology and complete results). Below we analyze performance patterns by task category and language, with full per-task results in Appendix B.
@@ 4.1 Performance Patterns by Task Category
Figure 2 shows human performance relative to model performance across all 26 task-language pairs. Each task shows human performance (point) positioned within the full spectrum from worst to best model performance (range bars). Humans consistently perform in the upper portion of model ranges, typically exceeding median model performance (61.5% of tasks) while rarely matching the best models (15.4% of tasks). Classification tasks show the strongest human performance, with humans outperforming all models in 3 of 7 tasks, while clustering and reranking reveal consistent gaps where humans fall short of top-performing models. Detailed gap analysis can be found in Appendix G.
Classification: Human performance averages 70.3, ranging from 45.8 on emotion classification ($\kappa=0.39$, fair agreement) to 95.0 on Arabic sentiment analysis. Models generally exceed human performance (best: 87.1), but humans outperform models on non-English sentiment analysis, particularly in Arabic (95.0 vs. 77.5) and Russian (92.5 vs. 81.2), likely benefiting from native cultural and linguistic understanding that current models fail to capture.
Clustering: Humans average 67.4 V-measure with extreme variation. Near-perfect performance on WikiCities (97.6, $\text{ARI}=0.91$) contrasts sharply with poor performance on ArXiv papers (49.2, $\text{ARI}=-0.001$). Models consistently outperform humans (best: 85.1%). The poor inter-annotator agreement on ArXiv indicates fundamental task ambiguity rather than human limitation.
Reranking: Humans achieve strong performance (87.2 average MAP) with high inter-annotator agreement ($\rho=0.64$-$0.85$), demonstrating intuitive relevance understanding. Models exceed human performance (best: 96.4), but the high human agreement suggests these tasks align well with human judgment and provide reliable evaluation targets.
STS: Humans average 83.2 Spearman correlation, with notable variation: STS12 achieves 91.2 while STS22-Russian drops to 58.7, likely reflecting dataset quality issues discussed in §4.3. Models achieve comparable performance (best: 88.5), with moderate inter-annotator agreement ($\rho=0.58$-$0.77$).